# SGDFuse: SAM-Guided Diffusion for High-Fidelity Infrared and Visible Image Fusion

### v2

@@ 4.2.3 Results and Analysis on the LLVIP Dataset
As shown in Table II, our proposed method attains the highest scores across all evaluation metrics
@@
except MI, where it ranks second, demonstrating strong competitiveness. These results highlight SGDFuse’s clear superiority in preserving information content, texture details, edge sharpness, and perceptual quality. Notably, leading performance in EN, SD, and SF indicates richer information, enhanced contrast, and improved clarity in the fused images. Furthermore, the superior $Q^{abf}$ and VIF scores confirm that SGDFuse achieves better perceptual consistency and edge alignment, closely aligning with human visual preferences.
The LLVIP dataset mainly contains nighttime scenes with challenges such as strong occlusion, weak textures, and extremely low illumination, which place higher demands on fusion robustness. Fig. 7 presents visual comparisons under these conditions. Methods like U2Fusion and TarDAL fail to enhance thermal saliency and suffer from blurred details and suppressed backgrounds. SHIP and SAGE improve thermal contrast but struggle with texture reconstruction and luminance balance. In contrast, our method preserves IR thermal responses while effectively integrating VIS textures and structures, producing results with sharper edges, consistent brightness, and stronger global contrast. Even under severe low-light conditions, SGDFuse demonstrates high stability and perceptual quality, confirming its robustness and practical applicability.
Figure 9: Comparison of different fusion methods in terms of parameter count and computational complexity.
@@ 4.2.4 Results and Analysis on the RoadScene Dataset
The RoadScene dataset includes complex traffic scenarios with varying illumination and diverse target distributions, imposing high demands on fusion for target detection and scene understanding. Fig. 8 shows visual comparisons. Methods such as U2Fusion and IGNet fail to retain both thermal cues and visual details, leading to blurred textures or weakened targets. Text-DiFuse and SAGE show limited target enhancement and detail clarity. MaeFuse preserves some textures but often produces dark results, reducing visual quality. In contrast, our method achieves superior brightness balance, structural clarity, and background suppression, demonstrating strong robustness and practical value under complex conditions.
4.3 Efficiency Comparison
To comprehensively compare with other state-of-the-art fusion methods, we evaluated the runtime efficiency of each approach in terms of computational cost and parameter scale. The detailed results are presented in Fig. 9. Computational cost is measured by the number of floating-point operations (FLOPs) required to process a single image, using the MSRS dataset for testing. TarDAL achieves higher efficiency due to its lightweight architecture with fewer parameters and lower computational cost; however, its fusion performance in terms of visual quality is inferior to our method. Diffusion-based methods introduce diffusion models, which significantly increase both parameter count and computational burden. SAGE incorporates SAM-generated segmentation masks as prompts to enhance semantic awareness in fusion but at the expense of increased computational complexity. Text-DiFuse employs a text-modulated diffusion model, improving flexibility and representational capacity in fusion, yet this also leads to higher computational overhead. In contrast, our method achieves a favorable balance by maintaining high-quality fusion results with moderate parameter size and computational cost, thus delivering better runtime efficiency. Overall, the analysis demonstrates that our approach strikes an effective trade-off between fusion performance and computational resource consumption, highlighting strong competitiveness and practical applicability.
@@ 4.4 Ablation Study
To thoroughly investigate the contribution of each component in the network architecture, a series of ablation studies were carefully designed and conducted, with systematic evaluations performed on the MSRS dataset.
4.4.1 Impact of SAM
The Segment Anything Module (SAM) offers high-quality semantic masks in stage II, guiding the diffusion model for structural refinement and semantic enhancement, thereby improving image integrity and consistency. To assess SAM’s effectiveness, we conducted ablation studies by removing
Table 3: Quantitative results of all fusion variants in the four ablation studies. Bold numbers indicate the best performance.
@@ IV. Impact of HFAH
Ours method 6.81 45.28 13.27 2.99 1.73 1.06 0.74
or modifying its components, designing three variants to analyze the impact of semantic guidance on fusion performance.
@@
(3). Variant (c): Removing the semantic mask of the visible image (w/o VIS SAM)
Contrary to the previous setting, this
@@
variant retains the infrared semantic mask while removing the visible light mask, thereby evaluating the importance of the rich texture and edge information from the visible images in detail preservation and target recognition.
As shown in Table III, variant (a) without SAM performs worst in all metrics, highlighting the importance of semantic priors for structural and semantic fidelity. Variants (b) and (c) show milder drops but remain inferior to the full model, confirming the complementary benefits of dual-modality semantic masks. These results validate SAM’s critical role in enhancing semantic representation and downstream task adaptability.
@@ 4.4.2 Analysis of the Impact of the Number of MSFEM and TB Modules
To explore the optimal stacking depth for feature extraction, we performed ablation experiments by varying the number of MSFEM and TB modules in Stage I from 2 to 4 to investigate the influence of stacking depth on fusion performance. As shown in Fig. 10, increasing the number of modules generally improves fusion quality by enhancing detail restoration and structural consistency. However, performance saturates or slightly declines beyond three layers due to potential redundancy and overfitting. Therefore, we adopt three MSFEM and three TB layers as the final configuration, striking a balance between accuracy and efficiency, and providing a strong feature foundation for stage II diffusion optimization.
@@ 4.4.3 Impact of Two-Stage Training
The proposed fusion framework adopts a two-stage training strategy: Stage I uses MSFEM and TB for modal feature extraction and preliminary fusion, while Stage II incorporates SAM and a conditional diffusion model for structural refinement under semantic guidance. To evaluate this design, we conducted ablation experiments comparing the full method with variants using only Stage I or Stage II. As shown in Table III, the Stage-I-only model (w/o Stage II) preserves basic structure but lacks detail and semantic consistency, while the Stage-II-only variant (w/o Stage I) yields blurred and distorted results due to missing structural priors. In contrast, the full two-stage strategy synergistically combines low-level structure with high-level semantics, enhancing alignment, detail preservation, and robustness in downstream tasks.
@@ 4.4.4 Impact of the Diffusion Process
As the core of the second stage, the diffusion model refines structure and enhances semantics via progressive denoising. To verify its effect, we conducted an ablation by removing the diffusion process (w/o Dif) while keeping the network unchanged. Table III shows that on MSRS, all metrics decline without diffusion, demonstrating its critical role in improving clarity, structural integrity, and fusion quality under semantic guidance. This confirms diffusion modeling’s unique advantage and necessity for high-fidelity, semantically consistent multimodal fusion.
@@ 4.4.5 Impact of HFAH
To enhance semantic guidance and structural restoration during denoising, we introduce the Hierarchical Feature Aggregation Head (HFAH) in the decoder of the diffusion model. HFAH aggregates multi-scale features via spatial attention to model edges and semantic regions more effectively. In ablation (w/o HFAH), removing hierarchical fusion and attention leads to noticeable drops in boundary fidelity and texture detail, as shown in Table III—especially in complex scenes. These results confirm that HFAH improves fine-grained structural recovery and semantic consistency.

### v4

@@ 4.2.3 Results and Analysis on the LLVIP Dataset
As shown in Table II, our proposed method attains the highest scores across all evaluation metrics except MI, where it ranks second, demonstrating strong competitiveness. These results highlight SGDFuse’s clear superiority in preserving information content, texture details, edge sharpness, and perceptual quality. Notably, leading performance in EN, SD, and SF indicates richer information, enhanced contrast, and improved clarity in the fused images. Furthermore, the superior $Q^{abf}$ and VIF scores confirm that SGDFuse achieves better perceptual consistency and edge alignment, closely aligning with human visual preferences.
The LLVIP dataset mainly contains nighttime scenes with challenges such as strong occlusion, weak textures, and extremely low illumination, which place higher demands on fusion robustness. Figure 8 presents visual comparisons under these conditions. Methods like U2Fusion and TarDAL fail to enhance thermal saliency and suffer from blurred details and suppressed backgrounds. SHIP and SAGE improve thermal contrast but struggle with texture reconstruction and luminance balance. In contrast, our method preserves IR thermal responses while effectively integrating VIS textures and structures, producing results with sharper edges, consistent brightness, and stronger global contrast. Even under severe low-light conditions, SGDFuse demonstrates high stability and perceptual quality, confirming its robustness and practical applicability.
@@ 4.2.4 Results and Analysis on the RoadScene Dataset
The RoadScene dataset includes complex traffic scenarios with varying illumination and diverse target distributions, imposing high demands on fusion for target detection and scene understanding. Figure 9 shows visual comparisons. Methods such as U2Fusion and IGNet fail to retain both thermal cues and visual details, leading to blurred textures or weakened targets. Text-DiFuse and SAGE show limited target enhancement and detail clarity. MaeFuse preserves some textures but often produces dark results, reducing visual quality. In contrast, our method achieves superior brightness balance, structural clarity, and background suppression, demonstrating strong robustness and practical value under complex conditions.
Table 3: Comparison of model parameters, latency, and FLOPs with other methods. Our latency is reported at our optimal T=60 steps. Latency for non-iterative methods is for a single forward pass.
Method Parameters(M) FLOPs(G) Latency(ms)
U2Fusion 0.66 517.34 104
TarDAL 0.30 221.33 41
PIAFusion 1.50 529.73 85
IGNet 7.87 51.49 98
CDDFuse 2.00 36.35 120
Diffusion 45.37 376.68 465
PFCFuse 0.64 302.75 35
SHIP 1.56 340.14 54
EMMA 1.52 106.41 64
Text-Difuse 121.51 243.68 350
DAFusion 0.28 8.24 132
MaeFuse 35.44 68.27 78
SAGE 0.14 52.06 25
Ours 47.16 220.74 59
4.3 Practical Efficiency Analysis
4.3.1 Computational Complexity and Inference Efficiency
To comprehensively evaluate the practical inference efficiency of the proposed framework, we conducted a rigorous comparison against other SOTA methods across three dimensions: parameter count (Parameters), computational cost (FLOPs), and average single-image inference latency (ms). As detailed in Table 3, the latency for non-iterative methods represents a single forward pass, while for iterative methods, we report the latency required to achieve optimal performance (T=60). When compared to similar diffusion-based models: Although our parameter count (47.16M) is comparable to that of Diffusion (45.37M), our SGDFuse (59ms) is significantly faster in terms of inference speed than both Diffusion and Text-Difuse. Furthermore, our practical latency (59ms) is also superior to many non-iterative methods, such as PIAFusion, and CDDFuse, and remains on the same order of magnitude as lightweight models like SAGE and TarDAL. In summary, this analysis demonstrates that our two-stage architecture (particularly the simplified diffusion task in Stage II) is computationally efficient. SGDFuse achieves a fast practical inference speed while maintaining a moderate computational cost (220.74G FLOPs). This proves that our method strikes an excellent trade-off between fusion performance and computational resource consumption, highlighting its strong competitiveness and practical applicability.
Figure 10: Comparative Analysis of VIF Performance versus Diffusion Steps on the MSRS Dataset.
Figure 11: Comparative Analysis of $Q^{abf}$ Performance versus Diffusion Steps on the MSRS Dataset.
4.3.2 performance metrics vs. number of diffusion steps
To further analyze the relationship between the number of diffusion steps and fusion performance, we plotted the ’quality-versus-speed trade-off curve’ for key performance metrics (VIF and $Q^{abf}$) as a function of sampling steps (T), and compared it against several representative SOTA methods, as shown in Figure 10 and Figure 11. Across all tested sampling steps, our method significantly and consistently outperforms other methods on both the VIF and $Q^{abf}$ metrics. Furthermore, it demonstrates excellent convergence efficiency. As illustrated in the figure, the model’s performance rapidly reaches its peak at T=60 steps. At this relatively low step count, our performance already comprehensively surpasses the baseline levels of all other SOTA methods. This perfectly demonstrates the excellent balance our method achieves between fusion performance and practical efficiency.
@@ 4.4 Ablation Study
To thoroughly investigate the contribution of each component in the network architecture, a series of ablation studies were carefully designed and conducted, with systematic evaluations performed on the MSRS dataset. Furthermore, to validate the generalizability of our core design choices, we conduct an additional ablation study on the challenging low-light LLVIP dataset, which represents a significantly different scenario.
Table 4: Quantitative results of key ablation variants on the MSRS dataset. Bold numbers indicate the best performance.
@@ IV. Impact of HFAH
Ours method 6.81 45.28 13.27 2.99 1.73 1.08 0.74
4.4.1 Impact of SAM
The Segment Anything Module (SAM) offers high-quality semantic masks in stage II, guiding the diffusion model for structural refinement and semantic enhancement, thereby improving image integrity and consistency. To assess SAM’s effectiveness, we conducted ablation studies by removing or modifying its components, designing three variants to analyze the impact of semantic guidance on fusion performance.
@@ 4.4.1 Impact of SAM
(3). Variant (c): Removing the semantic mask of the visible image (w/o VIS SAM) Contrary to the previous setting, this variant retains the infrared semantic mask while removing the visible light mask, thereby evaluating the importance of the rich texture and edge information from the visible images in detail preservation and target recognition.
As shown in Table 4, variant (a) without SAM performs worst in all metrics, highlighting the importance of semantic priors for structural and semantic fidelity. Variants (b) and (c) show milder drops but remain inferior to the full model, confirming the complementary benefits of dual-modality semantic masks. These results validate SAM’s critical role in enhancing semantic representation and downstream task adaptability.
@@ 4.4.2 Analysis of the Impact of the Number of MSFEM and TB Modules
To explore the optimal stacking depth for feature extraction, we performed ablation experiments by varying the number of MSFEM and TB modules in Stage I from 2 to 4 to investigate the influence of stacking depth on fusion performance. As shown in Figure 12, increasing the number of modules generally improves fusion quality by enhancing detail restoration and structural consistency. However, performance saturates or slightly declines beyond three layers due to potential redundancy and overfitting. Therefore, we adopt three MSFEM and three TB layers as the final configuration, striking a balance between accuracy and efficiency, and providing a strong feature foundation for stage II diffusion optimization.
@@ 4.4.3 Impact of Two-Stage Training
The proposed fusion framework adopts a two-stage training strategy: Stage I uses MSFEM and TB for modal feature extraction and preliminary fusion, while Stage II incorporates SAM and a conditional diffusion model for structural refinement under semantic guidance. To evaluate this design, we conducted ablation experiments comparing the full method with variants using only Stage I or Stage II. As shown in Table 4, the Stage-I-only model (w/o Stage II) preserves basic structure but lacks detail and semantic consistency, while the Stage-II-only variant (w/o Stage I) yields blurred and distorted results due to missing structural priors. In contrast, the full two-stage strategy synergistically combines low-level structure with high-level semantics, enhancing alignment, detail preservation, and robustness in downstream tasks.
@@ 4.4.4 Impact of the Diffusion Process
As the core of the second stage, the diffusion model refines structure and enhances semantics via progressive denoising. To verify its effect, we conducted an ablation by removing the diffusion process (w/o Dif) while keeping the network unchanged. Table 4 shows that on MSRS, all metrics decline without diffusion, demonstrating its critical role in improving clarity, structural integrity, and fusion quality under semantic guidance. This confirms diffusion modeling’s unique advantage and necessity for high-fidelity, semantically consistent multimodal fusion.
@@ 4.4.5 Impact of HFAH
To enhance semantic guidance and structural restoration during denoising, we introduce the Hierarchical Feature Aggregation Head (HFAH) in the decoder of the diffusion model. HFAH aggregates multi-scale features via spatial attention to model edges and semantic regions more effectively. In ablation (w/o HFAH), removing hierarchical fusion and attention leads to noticeable drops in boundary fidelity and texture detail, as shown in Table 4—especially in complex scenes. These results confirm that HFAH improves fine-grained structural recovery and semantic consistency.
Table 5: Quantitative results of key ablation variants on the LLVIP dataset. Bold numbers indicate the best performance.
LLVIP EN SD SF MI SCD VIF $Q^{abf}$
I. Impact of SAM
w/o SAM 6.87 42.26 15.69 2.52 1.43 0.83 0.63
w/o IR SAM 6.95 42.75 15.87 2.68 1.60 0.85 0.64
w/o VIS SAM 6.98 43.01 15.96 2.73 1.56 0.87 0.66
II. Impact of Two-Stage Training
w/o Stage I 6.91 42.94 15.87 2.60 1.55 0.93 0.64
w/o Stage II 6.82 42.79 15.39 2.59 1.53 0.91 0.62
III. Impact of the Diffusion Process
w/o Dif 6.99 43.07 15.98 2.65 1.54 0.92 0.65
IV. Impact of HFAH
w/o HFAF 7.06 43.19 16.11 2.72 1.61 0.96 0.69
Ours method 7.12 43.35 16.15 2.79 1.65 0.99 0.71
4.4.6 Generalizability Validation on LLVIP Dataset
To further validate the robustness and generalizability of our core design choices, we conducted an additional ablation study on the LLVIP dataset. This dataset, containing primarily nighttime scenes with weak textures and extremely low illumination , provides a significantly different and challenging scenario compared to the MSRS dataset. We re-evaluated our most critical components, with the results presented in Table 5. The results strongly mirror our findings from Table 4. The full model (Ours method) achieves the best performance. Notably, the w/o Stage II variant shows a severe degradation in all metrics, confirming that our diffusion-based refinement stage is essential for high-fidelity reconstruction across scenarios. Similarly, the w/o SAM and w/o HFAH variants also perform significantly worse than the full model, demonstrating that semantic guidance and hierarchical feature aggregation are robust and critical contributions. This consistent performance across both MSRS and LLVIP datasets confirms that the contributions of our core components are robust and generalizable.
Figure 13: Visualization comparison of the diffusion denoising process with (SAM) and without (w/o SAM) SAM guidance.
4.4.7 Semantically Guided Analysis of Diffusion Processes
To deeply analyze how SAM-based guidance influences the fusion process during diffusion, we visualized intermediate steps in the reconstruction of diffusion models. We compared the full model (SAM) with its variant without SAM. As shown in Figure 13, we observe that in the early stages ($T=60$, $T=45$), both trajectories are (as expected) dominated by noise. However, a significant divergence emerges at $T=30$: the full SAM model begins constructing coherent semantic structures for the left figure and distinct contours for the right vehicle, while the non-SAM model remains blurred noise at identical strides, failing entirely to capture these semantic targets. This semantic advantage persists through $T=15$ and into the final fusion stage. This visualization clearly demonstrates that our SAM-guided framework effectively steers the denoising trajectory toward semantically correct solutions from early stages, thereby revealing the model’s internal mechanisms with greater clarity.
4.5 Robustness and Generalizability Analysis
To comprehensively assess the stability and applicability of the proposed framework, this section conducts an in-depth analysis of two critical aspects: (1) the framework’s robustness to inaccuracies in semantic priors; and (2) its generalizability to different sources of semantic priors.
Table 6: Quantitative robustness analysis on the M3FD dataset using perturbed SAM masks. Bold numbers indicate the best performance.
Perturbation EN SD SF MI SCD VIF $Q^{abf}$
Eroded Mask 6.93 37.91 13.19 2.93 1.64 0.97 0.65
Dilated Mask 6.89 37.85 13.14 2.91 1.61 0.95 0.62
Original mask 6.99 38.02 13.27 2.99 1.69 1.01 0.69
Figure 14: Qualitative analysis of robustness to perturbed SAM masks.
4.5.1 Robustness Analysis to SAM Prior Inaccuracies
To address the concern that our framework’s performance may hinge on the accuracy of SAM-generated masks, we conducted a robustness experiment on the MSRS dataset. We simulated two common types of segmentation inaccuracies mentioned by the reviewer: ”incomplete masks” (False Negatives) and ”mis-segmented masks” (False Positives). To simulate incomplete masks and mis-segmented masks, we applied Morphological Erosion with a $7\times 7$ kernel to the original SAM masks ($M_{ir},M_{vis}$). We then fed these ”perturbed” masks into our frozen, pre-trained Stage II model to evaluate the impact on fusion quality.
This robustness analysis is visualized in Figure 14
Figure 15: Visual comparison of fusion results using different semantic guidance (Mask2Former, SegFormer, SAM) versus the unguided baseline (w/o SAM). The green boxes highlight the regions of interest (pedestrian targets) for clearer comparison of semantic preservation.
and quantitatively evaluated in Table 6. Figure 14 clearly demonstrates that the $7\times 7$ kernel causes significant perturbations along the mask boundary, yet the corresponding fusion results still exhibit high robustness. Despite minor visual changes compared to the original results, the model did not experience catastrophic failure nor produce the expected “false fusion emphasis” phenomenon. Key targets (e.g., figures) are preserved, and background textures remain stable. The quantitative analysis in Table 6 corroborates this conclusion. As anticipated, all metrics exhibit predictable declines after introducing the noise mask, confirming that semantic priors indeed guide the fusion process. The key finding, however, is that performance degradation is gradual rather than abrupt. For instance, the $Q^{abf}$ metric measuring structural fidelity decreased only from 0.69 to 0.65 (erosion) and 0.62 (dilatation). Similarly, the visual information fidelity (VIF) metric decreased only from 1.01 to 0.97 and 0.95. This experiment demonstrates through both visual and quantitative validation that while our framework significantly benefits from precise prior information, it is not overly fragile and maintains high robustness in the face of segmentation errors.
4.5.2 Generalizability Analysis to Alternative Semantic Priors
To verify whether our “semantic guidance” concept is inherently generalizable or merely fragilely dependent on a specific tool (i.e., SAM), we conducted a “substitute prior” experiment. Specifically, we replaced
Table 7: Generalizability Analysis of the Framework to Different Semantic Priors on the MSRS Dataset. Bold numbers indicate the best performance.
Method EN SD SF MI SCD VIF $Q^{abf}$
w/o SAM 6.43 43.27 11.98 2.54 1.48 0.87 0.62
w/ Mask2Former 6.72 44.95 12.97 2.90 1.67 1.03 0.71
w/ SegFormer 6.70 44.83 12.88 2.87 1.64 1.01 0.70
w/ SAM 6.81 45.28 13.27 2.99 1.73 1.08 0.74
SAM with Mask2Former and SegFormer, both pretrained on large-scale datasets such as COCO, as alternative sources of semantic priors. All other components of our model (including pretrained weights) were kept unchanged, and we re-evaluated the framework on the MSRS dataset. As shown in Table 7, using Mask2Former [37] or SegFormer [38] as guiding priors yielded significant performance improvements compared to the “no-guidance” baseline (w/o SAM). To visually demonstrate this point, we also conducted a qualitative comparison, as clearly shown in Figure 15. The ’w/o SAM’ (unguided) model fails to preserve critical target structures (such as pedestrians), whereas all three ’guided’ models (w/ Mask2Former, w/ SegFormer, and w/ SAM) successfully and robustly highlight the targets. This strongly demonstrates that the core idea of semantic guidance is indeed general and effective, rather than being narrowly tied to SAM. Moreover, the full version with SAM achieved the best overall performance across all key metrics. This indicates that, owing to its powerful open-set generalization capability, SAM provides the highest-quality semantic priors for this fusion task, thereby validating the rationale of adopting it as our preferred semantic guidance tool.
Table 8: Quantitative comparison on MRI-PET and MRI-SPECT datasets, with red representing the best results, blue representing the second best results.
Method MRI-PET MRI-SPECT
EN SD SF MI SCD VIF $Q^{abf}$ EN SD SF MI SCD VIF $Q^{abf}$
U2Fusion[1] 3.48 55.46 22.17 1.61 1.22 0.35 0.47 3.44 54.24 18.67 1.66 1.18 0.51 0.59
CDDFuse [35] 4.37 77.28 28.07 1.82 1.79 0.60 0.64 4.02 63.45 20.28 1.93 1.69 0.67 0.68
Dif-fusion[5] 4.34 74.12 27.13 1.81 1.66 0.59 0.64 3.95 63.31 20.10 1.83 1.64 0.63 0.67
SHIP [24] 5.34 67.96 26.33 1.84 1.53 0.60 0.65 4.79 52.20 19.77 1.88 0.67 0.60 0.68
Text-Difuse[11] 5.07 76.49 25.74 1.73 1.71 0.46 0.34 4.81 64.20 17.10 1.62 1.60 0.44 0.26
MaeFuse [6] 5.19 77.39 28.41 1.83 1.75 0.62 0.69 4.85 64.56 20.33 1.92 1.71 0.66 0.68
SAGE [12] 5.32 77.31 28.12 1.85 1.76 0.61 0.67 4.87 64.79 20.16 1.95 1.74 0.69 0.69
SGDFuse(Ours) 5.24 77.61 28.19 1.89 1.77 0.64 0.71 4.82 64.62 20.31 1.98 1.73 0.72 0.70
4.6 Validation on Medical Multimodal Fusion Domains
To validate the robustness and generalization capabilities of our proposed SGDFuse framework beyond the IVIF domain, we conducted additional cross-domain experiments. We evaluated the method on two challenging public medical image fusion datasets: MRI-PET and MRI-SPECT. Using the standard training-test split from the Harvard Medical website, we selected several representative state-of-the-art methods for comparative analysis.
Figure 16: Visual comparison of fusion results by our method against other representative state-of-the-art methods on the MRI-PET and MRI-SPECT datasets.
Quantitative results are presented in Table 8, while qualitative comparisons are shown in Figure 16. The results demonstrate our framework’s robust generalization capability. On the MRI-PET dataset, SGDFuse achieves state-of-the-art (SOTA) performance on key metrics including MI, VIF, and $Q^{abf}$. On the MRI-SPECT dataset, our model remains highly competitive. Figure 16 demonstrates that our original fusion method preserves finer structures and details in both PET and MRI images, showcasing its superior capability in retaining intricate details and enhancing overall image quality. This experiment confirms that our core “SAM-guided generation” paradigm is not confined to narrow IVIF solutions but represents a robust and generalizable framework that effectively adapts to other challenging domains, such as medical image fusion.